OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______ Numerical Answers Expected!

OpenStudy (aum):

f(n) = f(1) + f(2) + f(n - 1) f(3) = f(1) + f(2) + f(3-1) = f(1) + f(2) + f(2) = 2 + 3 + 3 = 8 f(4) = ? f(5) = ?

jhonyy9 (jhonyy9):

so you need to calcule first f(4) = f(1) +f(2) + f(4-1) = 2+3+8 =13 then f(5) = f(1) +f(2) +f(5-1) =2+3+f(4) = 5+13=18 do you understand it now ?

OpenStudy (anonymous):

So whats f(5)? @aum

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

@ganeshie8

jhonyy9 (jhonyy9):

f(5)=18